OpenStudy (nuccioreggie):

There are (37)2 ⋅ 30 cherries on a farm. What is the total number of cherries on the farm? 35 39 314 349

OpenStudy (nuccioreggie):

@Michele_Laino

OpenStudy (michele_laino):

hint: \[\Large \begin{gathered} {3^0} = 1 \hfill \\ {\left( {{3^7}} \right)^2} = {3^{7 \times 2}} = {3^{14}} \hfill \\ \end{gathered} \]

OpenStudy (nuccioreggie):

@Michele_Laino

OpenStudy (michele_laino):

please apply my hint

OpenStudy (nuccioreggie):

C

OpenStudy (michele_laino):

that's right!

OpenStudy (nuccioreggie):

C

OpenStudy (michele_laino):

yes! correct!

OpenStudy (nuccioreggie):

Ok thank you i didnt mean to type it twice i have more questions